OpenStudy (anonymous):

Solve for x and y by deriving an expression for y from the second equation, and substituting it back into the first equation. −x−3y=1 5x+3y=−8

OpenStudy (hoblos):

5x+3y=−8 ----> 3y = -5x -8 y = (-5/3)x -8/3 replace in equation 1 --> -x -(-5x -8) = 1 4x+8 =1 x = -7/4 y = (-5/3)x -8/3 = 35/12 - 8/3 = 3/12 = 1/4